An artist wants her painting to be in the shape of a golden rectangle. If the length of the painting is 36 inches, then what should be the width?
Answer by scott8148(6628)

a "golden" rectangle is one that when divided into a square and a rectangle (by a line across the width)
___ the resulting smaller rectangle is similar to the original (also "golden")
___ in other words, the ratio of the length to width is the same

in mathematical terms ___ L/W = W/(L-W)

"cross" multiplying __ L^2 - LW = W^2

36^2 - 36W = W^2 ___ 0 = W^2 + 36W - 1296

use quadratic formula to find W